(b) Calculate the missing frequency from the data if the median is 50.
CI 10-20 20-30 30-40 40-50 50-60 60-70
f 2 8 6 ? 15 10
1
Expert's answer
2019-10-25T11:15:22-0400
"\begin{array}{c:c:c}
 \begin{matrix}
 Class \\
 Interval\\ CI
\end{matrix}& \begin{matrix}
 Frequency \\
 f
\end{matrix} &\begin{matrix}
 Cumulative \\
 Frequency
\end{matrix} \\ \hline
 10-20 & 2 & 2 \\
 \hdashline
 20-30 & 8 & 10
\\
 \hdashline
 30-40 & 6 & 16\\
 \hdashline
 40-50 & f_4 & 16+f_4\\
 \hdashline
 50-60 & 15 & 31+f_4\\
 \hdashline
 60-70 & 10 & 41+f_4
\end{array}"


"Estimated\ Median=L+{{n \over 2}-B \over G}\times w"


where:

"L"  is the lower class boundary of the group containing the median

"n"  is the total number of values

"B"  is the cumulative frequency of the groups before the median group

"G"  is the frequency of the median group

"w" is the group width

We have that "n=2+8+6+f_4+15+10=41+f_4"


"Median=50"

The median lies in class interval "CI=50-60"

"L=50, n=41+f_4, B=16+f_4,G=15,w=10"


"50=50+{{41+f_4 \over 2}-(16+f_4) \over 15}\times 10""{41+f_4 \over 2}-(16+f_4) =0"


"41+f_4=32+2f_4"

"f_4=9"

The missing frequency is "9."